IR SPECTROSCOPY OF [Ag·(CO_{2})_{n}]^{−} CLUSTERS: IMPLICATIONS FOR REDUCTIVE ACTIVATION OF CO_{2}

\underline{BENJAMIN J. KNURR} and J. MATHIAS WEBER, JILA and Department of Chemistry and Biochemistry, University of Colorado, Boulder, CO 80309.

Reduction of CO_{2} is an essential step in its chemical conversion from a greenhouse gas to useable fuel stocks, but is energetically unfavorable. Association with anions has been shown to facilitate partial charge transfer to a target CO_{2} molecule, which could be a significant step towards CO_{2} recycling. However there is still much uncertainty in the role of solvent effects on these chemical processes. We present infrared spectra of [Ag·(CO_{2})_{n}]^{−} (n = 2−11) to elucidate the nature of the charge carrier in the cluster, the effects of solvation on the charge distribution and the amount of reductive activation of CO_{2} in the presence of a Ag anion. The structures of the [Ag·(CO_{2})_{n}]^{−} clusters are discussed in the framework of density functional theory. We compare and contrast the findings of [Ag·(CO_{2})_{n}]^{−} to those recently published on [Au·(CO_{2})_{n}]^{−} [1].

1. B.J. Knurr and J.M. Weber, J. Amer. Chem. Soc., 134 (2012) 18804-18808.